The sample space of a random experiment is fa, b, c, d, e, fI , and each outcome is equally likely. A random variable is defined as follows outcome abcde 0 03.9 7 715 Evaluate the cumulative distribution function of X at specified values. [Give exact answers in form of fraction.] F(0) F(3.9) F(7) F (15)
Answer #1

Given data,

Outcomes a b c d e f
X 0 0 3.9 7 7 15
probability 1/6 1/6 1/6 1/6 1/6 1/6

The cumulative distribution function of a real-valued random variable {displaystyle X} is the function given by


{displaystyle F_{X}(x)=operatorname {P} (Xleq x)}

Therefore from given data we calculate

1) F(0) = P(Xleq 0) = P(X=0) = 1/6

2) F(3.9) = P(Xleq 3.9) = P(X=0)+ P(X=3.9) = 2/6+1/6 =3/6

3)F(7) = P(X < 7) 11642|6 = 516 = P(X = 0)-P(X = 3.9)-P(X = 7) = 2/6+

15) = 2/64 1 /64 2/64 1/6-6/6

Therefore

x 0 3.9 7 15
p(x) 2/6 1/6 2/6 1/6
F(X) 2/6 3/6 5/6 6/6